Calculate the energy for one mole of photons (an einstein) for light absorbed at 7.00 × 102 nm.
Answer #1


Given:
lambda = 700 nm = 7.00*10^-7 m

Find energy of 1 photon first

use:
E = h*c/lambda
=(6.626*10^-34 J.s)*(3.0*10^8 m/s)/(7.00*10^-7 m)
= 2.84*10^-19 J
This is energy of 1 photon

Energy of 1 mol = energy of 1 photon * Avogadro's number
= 2.84*10^-19*6.022*10^23 J/mol
= 1.71*10^5 J/mol
= 1.71*10^2 KJ/mol
Answer: 171 KJ/mol